OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 48 when x = 8, find the value of f(x) when f(x) when x = 2 A.3 B.12 C.6 D.8

OpenStudy (anonymous):

2/8 = x/48 x = 12 The correct answer is B

OpenStudy (anonymous):

alrighty then thanks

OpenStudy (anonymous):

yw